# Latent Space Imaging

### v1

@@ 2.2 Physical Setup
To experimentally verify Latent Space Imaging, we constructed a proof-of-concept setup as depicted in Fig. 1a. In this setup, a scene is projected onto a Digital Micromirror Device (DMD TI DLP4500) using an objective lens. The image is then spatially modulated by the optimized patterns. The modulated output is captured by a relay lens, which focuses it onto a single-pixel detector (ThorLabs PDA100A2). This detector acts as an integrator, measuring a single value that represents the integrated intensity of the entire modulated image. A computer controls and synchronizes the operation of the DMD and the measurement signals, enabling the time-multiplexing of all measurements that compose the compressed signal.
Fig. 3c demonstrates the reconstruction quality of our experimental setup using a monitor as the source image. For such a task, we utilized the LSI trained to reconstruct images based on $256$ measurements. fine-tuned our model with 200 images from the FFHQ dataset, and used 100 unseen images from CelebAHQ to validate our results. This step accounts for non-modeled noise and sensor non-linearities. More details on the supplementary material.
Refer to caption a c b
Figure 3: (a) Shows the histogram of pixel occupancy, comparing scenarios where the energy loss function is and is not utilized. The use of energy modeling enhances the dynamic range of the measurements captured by the single sensor. (b) Displays representative patterns after using the energy loss function. (c) Demonstrates real-world captures using the setup depicted in Fig. 1c. These images are acquired with only 256 measurements, which constitutes merely approximately $0.4\%$ of the total pixel count, based on an image resolution of $256\times 256$.
3 Discussion
Compared to other imaging techniques, Latent Space Imaging (LSI) offers substantial benefits in domain specific scenarios with tight requirements for hardware complexity, memory and bandwidth requirements, or frame rate. As Fig. 2b illustrates, LSI is capable of keeping details even in a compression ratio 1:1,000, represented by the 64 out of 65536 measurements case, were the overall facial structure, expression, eyes and skin are well preserved.
@@ 3 Discussion
The LSI reconstruction power inherits the advantages of a well-trained generative model specialized in a specific domain. In this approach, we leverage a pre-trained StyleGAN2, optimized to cover the entire distribution of the dataset. Consequently, its latent space learns to represent the most meaningful and essential information within this domain. This capability enables high-quality measurements with high compression ratios. However, it also inherently possesses the limitations of the underlying generative model and the inversion network.
Therefore, the model may fail to reconstruct very specific features not present during training, such as hats, glasses, makeup, or other distinct image details, and may struggle to generalize to a multi-domain approach. Additional analysis about challenging scenarios and how different generative models can be utilized on out framework can be found at supplementary materials.
3.1 Alternative Physical Implementations
@@
However, SPI also has several downsides that limit its practicality for real imaging systems, including its form factor, and, most notably, its difficulty in dealing with moving scenes. However, the recent efforts in optical computing [28] have resulted in a range of alternative options for implementing the required optical encoding layer. For example diffractive optical elements have been used to implement both convolutional [29] and fully connected linear layers [30], and meta-surfaces have been demonstrated for shift variant convolution kernels [31]. Due to the linear nature of light, all these methods are fundamentally limited to linear operators. However, as we show with the SPI setup, linear optical encoders are sufficient for achieving high compression rates when combined with a small digital encoder. As such, any of the recently proposed optical compute frameworks could be used instead of the SPI framework, although typically this would require “freezing” the optical encoder into hardware. Nonetheless, for special purpose imaging systems, such an approach could present a more practical avenue than the SPI framework.
3.2 Relationship to Biological Vision
As mentioned in the introduction, LSI is inspired by biological vision, in particular the human visual system (HVS). Specifically, in the HVS, photoreceptors sense the spatial distribution of light, retinal ganglion cells (RGCs) encode these spatial distributions into a compressed latent space transmitted over the optic nerve, and finally the visual cortex decodes and processes the information. LSI can be seen as analogous to this process: in our single pixel camera setup, the pixels of the spatial light modulator take on the role of the photoreceptors – they define the spatial resolution of the sensed image. The optical codes shown on the SLM, combined with the digital encoder take on the role of the RGCs, and both encode and compress this information, while the StyleGAN takes on the role of the visual cortex.
The encoding module is therefore composed of a combination of a (linear) optical computing part and a non-linear digital encoder. The digital encoder is needed since optical computing [28] is in practice limited to linear operators. Nonetheless, our work shows that even with linear optical computing, data reduction in the range of 100:1 to 1,000:1 is feasible even for the challenging task of full image restoration. Easier tasks such as object detection or classification will likely be possible with substantially higher compression ratios.
4 Future Potentials of LSI and Conclusion
Refer to caption a b c d
Figure 4: Visualization of image and latent space representations using UMAP. (a) UMAP projection of images reconstructed by our pipeline, using images from the test set of CelebA-HQ, colored by gender and highlighting the image 172631 and its neighborhood. (b) UMAP projection of the corresponding latent space representations, colored by the same categories. (c) Visualization of the chosen image in the image space with its four closest neighbors. (d) The same example image with its four closest neighbors in the latent space, highlighting stronger semantic similarities among neighbors. We fit the UMAP for 5 nearest neighbors for each point, with a small minimum distance of 0.001 to ensure that points are tightly packed, preserving local relationships and creating distinct clusters.
Latent Space Imaging provides a general framework for imaging within a compressed environment, but its utility is not limited to this application alone. This technique opens up new avenues for exploring deeper interactions and potential innovations in image processing and analysis.
Previous studies have delved into the semantics within GAN models, uncovering boundaries in the latent space that distinguish various image attributes [32, 33, 34]. Direct visualization of these models’ latent spaces allows for a natural observation of these structures, offering significant advantages for various applications without the need for additional complex deep learning frameworks to extract these semantics.
To further investigate this, we analyzed the latent space representation L of each image in the CelebA-HQ [35] test set using our pipeline, along with images I’ reconstructed by StyleGAN2. We performed UMAP [36] feature reduction to visualize them in a 2D map. Fig.4a shows the image representation, while Fig.4b illustrates the latent space representation.
The latent representation forms clusters based on image features. For instance, we highlighted males, females, and one specific image along with its neighborhood in different colors. Males tend to cluster on one side, separated from females, and image 172631 clusters close to images with similar characteristics (e.g., hair color, age, glasses, skin tone). This clear semantic division is not observable in the image-based representation. To illustrate this, we examined the four closest neighbors in each representation. It is evident that the neighbors in the latent space share more semantic similarities (see Fig. 4c,d).
This analysis not only underscores that latent space representations enable a novel compressed imaging process, but also demonstrates their potential as powerful semantic extractors. Such capabilities suggest that exploring these representations could yield significant insights and advancements in the field of generative models, potentially leading to more effective and nuanced image processing techniques in the future.
Besides a semantic analysis, the LSI framework here leverages a single-pixel approach. This implementation can be engineered for high frame rate due to low bandwidth requirements, and utilizing state-of-the-art DMDs would provide an out-of-box improvement for the already proposed setup. As discussed, Latent Space Imaging (LSI) could also be effectively implemented using a variety of other optical computing paradigms, such as diffractive optical elements and meta-surfaces (see Sec.3.1). These alternatives to the traditional Single Pixel Imaging (SPI) framework can manage optical encoding with high efficiency. Furthermore, enhancing the digital decoder through research into different generative model implementations holds promise for significantly improving both the versatility and efficacy of LSI systems. This approach could lead to more practical and adaptable imaging solutions, particularly suited for dynamic environments.
5 Methods
We utilized PyTorch to design and validate the LSI framework. All training details, hyperparameters and $\mathcal{D}_{\theta}$ architecture detailed description and discussion can be found in the supplementary material.
5.1 Optical Encoder Optimization
In the O optimization process, we employ the straight-through estimator (STE) [27], a technique also favored (VQ-VAEs) [37, 38, 39]. During the forward pass, values are quantized and constrained before computing J. However, during back-propagation, gradients are allowed to flow through the non-quantized version, facilitating weight updates. Importantly, we ensure that the entries of O remain positive and are bounded between $[0,1]$. Although using complementary patterns [40] allows for the inclusion of negative values, empirical experiments have shown that this approach does not yield significant improvements and necessitates double the measurements.
Initially, masks are generated from a uniform distribution and are then binarized through a quantization process. This process results in binary patterns that achieve an even total counting of 0s and 1s, theoretically maximizing bit entropy and steering the system towards an optimal solution [41]. This distribution presumes the sensor is sensitive enough to discern very fine differences, a demanding prerequisite as each pattern will have similar total intensity. This issue is addressed using the energy loss mechanism, which was introduced in Section 2.2 and is further detailed in the supplementary material.
5.2 Digital Encoder Optimization
Inspired by inversion networks, our approach tailors the digital encoder specifically for our unique imaging task. Unlike traditional inversion processes used for image editing, our encoder is designed for an optically compressed inversion, aligning the characteristics of the latent vector with the optical system’s constraints. This method allows our encoder to generalize across any image within the domain, without the need for training on specific data points.
In the broader context of the traditional inversion process, a latent array is retrieved from a pre-trained generative model to accurately reproduce the original image. This enables not only spatial adjustments, such as translation and rotation, but also semantic modifications including age, hair color, and expression changes. The inversion can be achieved through various methods: it may involve a full optimization for each data point [42, 21, 43], the use of an encoder trained to produce latent representations in a zero-shot manner [44, 24], or a combination of these strategies [45, 46, 47].
5.3 Datasets
To effectively exploit the latent space of generative models, we have narrowed our focus to the domain of human facial recognition. Specifically, we aim to retrieve identities from pictures of human faces. For this purpose, we utilized two datasets: FFHQ [20] and CelebAHQ [35]. The FFHQ dataset was exclusively employed for training, while CelebAHQ was used both for training and testing—comprising 28,000 images for training and 2,000 for testing. Combined with FFHQ, this resulted in a total of 98,000 face images used for training. This approach ensures a robust dataset that enhances the reliability and effectiveness of our model in capturing and reconstructing facial identities.
Both datasets contain background information that could pose a challenge due to the already compressed measurements. To address this, we further refine the domain by applying Segment Anything (SAM) [48] to remove background elements and prioritize identity retrieval. Moreover, the versatility of LSI allows it to be easily extended to other specific domains by adopting generative models trained on different datasets.
6 Code Availability
The code will be made available together with all checkpoints to ensure reproducibility.
7 Data Availability
The CelebA-HQ dataset is accessible on GitHub at https://github.com/tkarras/progressive_growing_of_gans. The FFHQ dataset can be found at https://github.com/NVlabs/ffhq-dataset. The AFHQ can be obtained at https://github.com/clovaai/stargan-v2/tree/master.
Supplementary Material
S1 Digital Encoder Architecture
Refer to caption a b

### v2

@@ 5 Hardware Prototype
A schematic of our setup is depicted in Fig. 1. In this system, a scene is projected onto a Digital Micromirror Device (DMD TI DLP4500) using an objective lens. The image is then spatially modulated by the optimized patterns. The modulated output is captured by a relay lens, which focuses it onto a single-pixel detector (ThorLabs PDA100A2). This detector acts as an integrator, measuring a single value that represents the integrated intensity of the entire modulated image. A computer controls and synchronizes the operation of the DMD and the measurement signals, enabling the time-multiplexing of all measurements that compose the compressed signal.
In the single-pixel camera implementation of the LSI framework, O spatially modulates the incoming image I through matrix multiplication (denoted by $\cdot$ in Eq. 1). Each row of the optical encoder corresponds to one learned mask, and each column represents one measurement. Each mask corresponds to a single pixel, where its forms are learned for a specific domain/imaging purpose. Having the hardware in the loop during training involves maintaining O’s entries as positive, bounded between [0, 1], and quantized. In order to maintain the proposed setup as minimal as possible, we utilize binary optical masks - corresponding to quantizing its values to 0 or 1. Such decision enables reaching high frame-rate on light modulators like Digital Mirror Devices (DMDs).
With this framework established, we further improve the system’s light efficiency and sensitivity to subtle variations by enforcing a $1\%$ intensity difference between each optical mask. This distinction is achieved by introducing an energy efficiency loss in the later stages of training, fine-tuning the encoding process to ensure each pattern is uniquely identifiable, thereby significantly enhancing the system’s ability to effectively discriminate between different patterns. Fig. 2 illustrates that the masks result in patterns highlighting the system’s focus on facial landmarks for specific domain encoding. Please refer to the Supplement for details.
Figure 2: Displays optimized pixel forms for 1:1024 compression ratio.
6 Training Details
Figure 3: Illustration of reconstructions achieved through the LSI pipeline with bounded and quantized pixels, using compression ratios from 1:128 to 1:4096 in the simulation setting.
Optical Encoder Optimization In optimizing O, we use the straight-through estimator (STE) [23], as in VQ-VAEs [62, 50, 43]. Forward passes quantize and constrain O values, while back-propagation allows gradients to flow through the non-quantized version, facilitating effective weight updates.
Datasets We utilize FFHQ [28] and CelebAHQ [27], reserving 2000 images from CelebAHQ for testing. Only CelebAHQ is used for downstream applications, where Segment Anything (SAM) [36] focuses on faces, disregarding background content. The Supplement includes additional evaluations on the AFHQ dataset [8].
Loss Functions Alongside the latent 3 and energy 5 losses, we employ the identity loss ($\mathcal{L}_{id}$), calculating cosine distance between feature maps from ArcFace [13]. We also apply the $\ell_{2}$ norm ($\mathcal{L}_{l2}$) and DINO/LPIPs features as perceptual loss ($\mathcal{L}_{p}$[79, 76].
The Supplement provides full details on hyperparameters, optimizers, loss functions, as well as downstream application losses and additional training technicalities.
7 Simulation Results
In our experiments we evaluate the LSI pipeline with simulated results as well as real-world measurements. We focus on the face domain understanding how the minimal imaging performs on essential facial tasks like reconstruction, identity recognition, attribute classification, segmentation, and landmark detection, analyzing the boundaries for each one of such tasks.
For quantitative results, we select two metrics for face identification: VGGFace [47] and Dlib [32]. In both cases, pre-trained face recognition models are used to extract features from the original and reconstructed images. Cosine similarity is then computed to verify identity; a similarity score below the thresholds of 0.68 and 0.07, respectively, indicates a match. These thresholds are sourced from the DeepFace [54] library. Additionally, FID [20] is also computed to compare the distribution of features between the reconstructed and the original set. Traditional metrics, such as PSNR and SSIM, are useful for evaluating pixel-level accuracy but often fall short in fully capturing human visual perception [76, 38, 2]. Further assessment and an in-depth discussions of these metrics are provided in the Supplement.
[…]
1:128 91.97% 92.74% 27.38 89.07% 70.00% 1.48
1:256 90.98% 92.68% 26.62 89.15% 70.94% 1.43
1:512 89.61% 91.67% 28.66 89.20% 70.25% 1.48
1:1024 81.12% 87.44% 28.79 88.74% 69.18% 1.52
1:2048 54.72% 77.77% 35.89 88.06% 65.81% 1.67
1:4096 27.22% 59.21% 46.18 86.44% 60.63% 2.01
1:8192 N/A N/A N/A 83.81% 53.77% 2.18
1:16384 N/A N/A N/A 81.75% 46.36% 2.47
Table 1: Latent Space Imaging is quantitatively evaluated for simulated results across multiple downstream tasks under highly compressed settings. For the image reconstruction task, validation is performed using both the VGGFace [12] and DLib [32] image recognition pipelines. Performance metrics include Fréchet Inception Distance (FID), classification accuracy, F1-mean score for segmentation tasks, and Normalized Mean Error (NME), normalized by the diagonal of the face bounding box. In cases of extreme compression ratios, while recognition metrics and reconstruction may not be applicable (N/A), downstream applications remain feasible.
7.1 Image Reconstruction
Full image reconstruction is the most demanding task for any reduced sampling scheme, especially in the case of human faces, where the human visual system is extremely sensitive to minute differences. Fig. 3 demonstrates the robust identity preservation capabilities of our model when reconstructing images at a resolution of $256\times 256$ across a range of measurement levels. Latent Space Imaging effectively captures essential features such as eye color, expression, facial shape, and skin tone with less than $0.4\%$ of the original pixel count. This capability is achieved by designing O to selectively retrieve relevant features from the target domain, enabling focused extraction of critical characteristics for facial identification while preserving landmarks embedded in the underlying distribution learned by the pre-trained generative model $\mathcal{G}$. The model $\mathcal{G}$ was trained exclusively on face images (FFHQ 256x256), which is the same training data used for all comparison methods. This targeted feature retrieval, along with compressing the target space into a flat latent structure, enables the reconstruction process not only to interpolate under-sampled data but also to generate outputs closely resembling the original, detailed facial images.
The compression ratios shown in Fig. 3 correspond to measurement levels of $512$, $256$, $128$, $64$, $32$, and $16$, with quantitative assessments provided in Table 1. Crucially, only $64$ measurements are already sufficient for accurate recognition with ArcFace [12] and Dlib [32], with accuracy declining significantly below $32$ samples. However, even for these extreme compression ratios, critical facial features—such as facial structure, gaze direction, and expression—are remarkably preserved. This underscores the advantage of targeting the latent space of generative models over achieving pixel-perfect accuracy. Beyond visualization, the linear boundaries within the latent space offer significant benefits for downstream applications, enabling effective feature recognition with minimal adjustments.
8 Prototype Results
Figure 4: Face reconstructions from the experimental setup across different compression ratios.
Real world experiments are conducted using our prototype and the pipeline described in Sec. 5. A set of 100 face images is randomly selected from CelebAHQ test set and displayed on a high-dynamic-range Eizo CG3145 monitor, then captured by the single-pixel setup. Fig. 4 illustrates the reconstruction results across three different compression ratios. These same measurements are further utilized by the linear projectors to produce attributes classification, segmentation masks and landmark points for compression ratios of 1:512, 1:2048 and 1:8192 (see Figs. 57 and 6), corresponding to $128$ and $32$ scalars, respectively, thereby validating the feasibility of the LSI framework.
In the following, we demonstrate how our sensing matrix, optimized for the latent space, effectively supports a wide range of downstream computer vision tasks without requiring full image reconstruction. All results presented are based on real-world experiments conducted using the prototype described in Sec. 5.
Each application is evaluated on the CelebAHQ [27] dataset, following the architectures and feature projections detailed in Sec. 4. For attribute classification, we use the original 40 labels per image as ground truth. In landmark detection, we leverage FaRL [78] as a state-of-the-art model to compute 68 landmark points, using the original image I as the reference. For segmentation tasks, we empirically found that generating masks using reconstructed faces I’ with FaRL as ground truth achieves superior accuracy, given the spatial precision needed for effective segmentation.
8.1 Attributes
The classification accuracy of 40 binary attributes is evaluated across the full range of compression ratios. Notably, we achieve $80\%$ precision even at a 1:16384 compression ratio, which corresponds to only 4 scalar measurements. Fig. 5 visualizes the sigmoid-transformed logits after linearly projecting the latent space, $P_{A}(\textbf{L})$. Despite the decline in similarity observed in Table 1 and Fig. 3, the selected attribute ratios remain close to the original, offering an accurate representation of the original face.
Figure 5: The figure illustrates the sigmoid-transformed logits for compression ratios of 1:256 and 1:2048, overlaid for comparison. These correspond to 256 and 32 scalars, respectively.
8.2 Facial Landmarks
Figure 6: Overlay of predicted and ground-truth positions for 68 facial landmarks across different compression ratios.
Fig. 6 highlights landmark detection acquired by the linear projection $P_{L}(\textbf{L})$ on the same two images across various compression levels, demonstrating high accuracy even in extremely compressed scenarios, such as 1:8192 (8 scalars), compared with the pseudo ground-truth ($gt$) generated by FaRL [78]. The third row shows a non-frontal view of a face, yet the predictions remain close to the reference, indicating robustness.
The quantitative evaluation in Table 1 uses the Normalized Mean Error (NME) metric (Eq. 4), scaled by $\alpha=100$ for visualization purposes. NME is calculated as the average Euclidean distance between predicted and ground-truth landmarks, normalized by the diagonal of the face bounding box ($d^{\text{dg}}$) to account for variations in face size.
$\text{NME}_{\text{dg}}=\frac{\alpha}{N}\sum_{i=1}^{N}\frac{\|P_{L}(\textbf{L})_{i}-{gt}_{i}\|_{2}}{d^{\text{dg}}_{i}}.$ (4)
8.3 Face Segmentation
The segmentation task is performed using the FaRL model, trained on the CelebAHQ dataset, which includes 18 classes. During our linear projection, $P_{S}$, we aim to learn the entirety of these classes. However, due to the limitations of our generative model, reproducing high-frequency details and unique objects are challenging; for instance, accessories like earrings are not well segmented. Despite this, for general facial features such as nose, hair, face, mouth, ears, and eyebrows, the linear projection of $\mathcal{G}$ internal features performs well, even under highly compressive settings.
Fig. 7 illustrates segmentation results across a range of different hair styles and different facial hair, showing that the segmentation masks remain consistent despite a four times increase in compression ratio. Additionally, Table 1 highlights the stability of the F1 score across different compression ratios, demonstrating reliable segmentation performance.
Figure 7: Illustration of face segmentation for two compression levels (1:512 and 1:2048) as well as the pseudo ground-truth for extract from the original image utilizing the FaRL model.
Extended results and discussions for all compression ratios, downstream tasks, evaluation metrics, detailed descriptions of the physical setup are provided in the Supplement.
9 Comparison with Alternative Approaches
Although our approach diverges from current strategies in compressed imaging, we selected notable models for comparison. Specifically, we included representative models from block-based sensing matrices [58, 17], unfolding networks [64], traditional single-pixel methods with deep learning-based reconstruction [16, 63], Fourier pattern single-pixel methods [77], and optimization-based methods using generative models as priors [24]. Table 2 and Fig. 8 illustrate the comparison quantitative and qualitative simulated results for deep learning single-pixel using Fourier features [77] and unfolding networks [64]. For fair comparison, the sensing matrix for the unfolding network is 1 bit and 8 bit quantized which is required for a plausible physical implementation. Quantization training follows the same strategy adopted by LSI. Detailed methodologies and all mentioned comparisons are elaborated in the Supplement.
Fig. 8 highlights the details of our model which contrasts with traditional compressive sensing strategies, where limited measurements lead to reconstruction algorithms over-smoothing the images, resulting in the loss of critical identity features such as the shape of eyes, facial hair, and other distinguishing landmarks. Table 2 also highlights that even with better quantization deep unfolding methods still outperformed by LSI.
Figure 8: Qualitative comparison between different methods.
[…]
FSI-DL 3.30% 13.15% 123.5
SAUNet 17.84% 50.12% 75.53
SAUNet 39.47% 65.97% 65.79
LSI 90.98% 92.68% 26.62
Table 2: Quantitative comparison between different compressed imaging methods for a 1:256 ratio. SAUNet represents a version retrained on 8 bits quantized sensing matrix, while the SAUNet was 1 bit quantized, similar to LSI and FSI-DL.
10 Discussion and Conclusion
Compared to other imaging techniques, Latent Space Imaging offers substantial benefits in domain-specific scenarios with tight requirements for hardware complexity, memory and bandwidth requirements, or frame rate. As mentioned in the introduction, LSI is strongly inspired by biological vision, particularly the human visual system. The relationship between LSI and biological vision is discussed in more detail in the Supplement.
@@ 10 Discussion and Conclusion
LSI inherits the advantages of a well-trained generative model specialized in a specific domain. Consequently, its encoding learns to represent the most meaningful and essential information within this domain. This capability enables high-quality measurements with high compression ratios. However, it also inherently possesses the limitations of the underlying generative model and the inversion network. Therefore, the model may fail to reconstruct very specific features not present during training, such as hats, glasses, makeup, or other distinct image details, and may struggle to generalize to a multi-domain approach.
Nonetheless, we believe that LSI is a promising new paradigm for domain-specific imaging systems, including privacy-preserving cameras that nonetheless excel at a specific imaging task. LSI is also a promising avenue for imaging systems ultra-low hardware requirements (both in terms of optical/electronic complexity and compute power), and/or ultra-high frame rate imaging. These applications will require alternative hardware implementations compared to our single-pixel camera prototype, which was chosen primarily due to easy reconfiguarbility. Several concepts for such alternative hardware designs are discussed in the Supplement.
Supplementary Material
[…]
Matheus Souza        Yidan Zheng        Kaizhang Kang        Yogeshwar Nath Mishra
                   Qiang Fu              Wolfgang Heidrich
KAUST
This supplementary document consists of three parts: I.) A more in-depth discussion of fundamental principles of Latent Space Imaging, including a discussion of the relationship to biological vision, and of alternative optical encoding schemes. II.) Additional implementation details of the current system, including the hardware prototype and training details. III.) Additional experimental results and ablation studies.
A Part I: Foundations of LSI
A.1 Relationship to Biological Vision
As mentioned in the main text, LSI is strongly inspired by biological vision, in particular the human visual system (HVS). Specifically, in the HVS, photoreceptors sense the spatial distribution of light, and retinal ganglion cells (RGCs) encode these spatial distributions into a compressed latent space transmitted over the optic nerve. Finally the visual cortex decodes and processes the information. The compression ratio between the number of photoreceptors and the number of axions in the optic nerve is about 100:1, and is enabled by the inherently non-linear processing of the RGCs.
LSI can be seen as analogous to this process: in our single pixel camera setup, the pixels of the spatial light modulator take on the role of the photoreceptors – they define the spatial resolution of the sensed image. The optical codes shown on the SLM, combined with the digital encoder take on the role of the RGCs, and both encode and compress this information, while the StyleGANXL takes on the role of the visual cortex.
The encoding module is therefore composed of a combination of a (linear) optical computing part and a non-linear digital encoder. The digital encoder is needed since optical computing [68] is in practice limited to linear operators. Critically, the most compressed representation is the intermediate latent space between the optical and the digital encoder, which enables highly compressed, low-bandwidth sampling hardware. We show that, with this two stage encoding, very high compression ratios can be achieved for the specific domain of facial images: ratios of 1:100 to 1:1000 for full image reconstruction, and even higher ratios for simpler tasks such as landmark detection or segmentation.
The principle of LSI can be applied to the latent space of any generative model, however the achievable compression ratios will depend on the complexity of the domain. To match the performance of the HVS on general vision tasks with a compression ration of 1:1000, we will need both a general AI generative model as well as a way to optically implement non-linear encodings.
A.2 Alternative Physical Implementations
@@
However, SPI also has several downsides that limit its practicality for real-world imaging systems, including its form factor, and, most notably, its difficulty in dealing with moving scenes. However, the recent efforts in optical computing [68] have resulted in a range of alternative options for implementing the required linear optical encoding layer.
@@
For example diffractive optical elements have been used to implement both convolutional [6] and fully connected linear layers [40], and meta-surfaces have been demonstrated for shift variant convolution kernels [67]. Due to the linear nature of light, all these methods are currently limited to linear operators. However, as we show with the SPI setup, linear optical encoders are sufficient for achieving high compression rates when combined with a small digital encoder. As such, any of the recently proposed optical compute frameworks could be used instead of the SPI framework, although typically this would require “freezing” the optical encoder into hardware. Nonetheless, for special purpose imaging systems, such an approach presents a practical avenue for developing real-world, snapshot-capable LSI systems.
In the longer term, LSI may be able to benefit from ongoing research efforts to develop non-linear optical computing hardware, for example based on non-linear optical materials, quantum effects, polarization, or other non-linear effects. Eventually, this may enable more powerful optical encoders, thereby reducing or even eliminating the need for the digital encoder module.
B Part II: Implementation Details
B.1 Digital Encoder Architecture